OpenStudy (anonymous):

A circle is measured by a straight line across its widest part. This line is the __________ of the circle. A. area B. volume C. diameter D. radius

OpenStudy (anonymous):

Answer: C diameter  d is diameter r is radius

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

What is the surface area of a circle that has a diameter of 14 inches? A. 49.86 in2 B. 153.86 in2 C. 196.86 in2 D. 616.86 in2

OpenStudy (anonymous):

Area=radius^2*pi =49*pi =49*3.14 =153.86 So B

OpenStudy (anonymous):

When the numerator of a fraction is smaller than the denominator, it is a(n) __________ fraction. A. improper B. reduced C. mixed D. proper

OpenStudy (anonymous):

Question 1 of 15 6.66 Points If John had 6 apples and Nicole had twice as many, how many apples did the two have in total? A. 6 B. 24 C. 18 D. 12 Reset Selection Mark for Review What's This? Question 2 of 15 6.66 Points The fraction 7/4 is known as a(n) __________ fraction. A. reducing B. improper C. proper D. whole Reset Selection Mark for Review What's This? Question 3 of 15 6.66 Points The decimal number 0.75 means: A. there are seven-and-a-half parts of a whole unit. B. seventy-five tenths of a whole unit. C. that a whole unit consists of seventy-five parts. D. there are seventy-five parts of a whole unit that has 100 parts. Reset Selection Mark for Review What's This? Question 4 of 15 6.66 Points A circle is measured by a straight line across its widest part. This line is the __________ of the circle. A. area B. volume C. diameter D. radius Reset Selection Mark for Review What's This? Question 5 of 15 6.66 Points What is the surface area of a circle that has a diameter of 14 inches? A. 49.86 in2 B. 153.86 in2 C. 196.86 in2 D. 616.86 in2 Reset Selection Mark for Review What's This? Question 6 of 15 6.76 Points When the numerator of a fraction is smaller than the denominator, it is a(n) __________ fraction. A. improper B. reduced C. mixed D. proper Reset Selection Mark for Review What's This? Question 7 of 15 6.66 Points Subtract: 1586 - 987 = ? A. 559 B. 699 C. 499 D. 599 Reset Selection Mark for Review What's This? Question 8 of 15 6.66 Points The fraction 25/100 can be reduced to 5/20, a fraction having the same value, by dividing both the numbers 25 and 100 by: A. 25 B. 12 C. 10 D. 5 Reset Selection Mark for Review What's This? Question 9 of 15 6.66 Points Divide: 100/22 = A. 4.345 B. 5.011 C. 4.545 D. 4.98 Reset Selection Mark for Review What's This? Question 10 of 15 6.66 Points Convert the mixed number, 1 3/4 , to a fraction. Reduce fraction if necessary. A. 4/12 B. 4/7 C. 12/4 D. 7/4 Reset Selection Mark for Review What's This? Question 11 of 15 6.66 Points The surface area of any shape is expressed in: A. length and width. B. square units. C. feet and inches. D. cubic units. Reset Selection Mark for Review What's This? Question 12 of 15 6.66 Points Reduce the fraction 20/100. A. 4/20 B. 2/20 C. 1/5 D. 5 Reset Selection Mark for Review What's This? Question 13 of 15 6.66 Points The number “Pi” (∏) is often rounded to what number? A. 3.15 B. 3.14 C. 3.01 D. 3.11 Reset Selection Mark for Review What's This? Question 14 of 15 6.66 Points Decimal numbers with digits both before and after the decimal point, such as 16.75, are referred to as: A. the remainder. B. decimal point numbers. C. mixed decimal numbers. D. the quotient. Reset Selection Mark for Review What's This? Question 15 of 15 6.66 Points Solve: 1000 + 69 - 55 = A. 986 B. 968 C. 1014 D. 1041 Reset Selection